# Worth Their Weight: Randomized and Regularized Block Kaczmarz Algorithms without Preprocessing

### v1

@@ Abstract
Due to the ever growing amounts of data leveraged for machine learning and scientific computing, it is increasingly important to develop algorithms that sample only a small portion of the data at a time. In the case of linear least-squares, the randomized block Kaczmarz method (RBK) is an appealing example of such an algorithm, but its convergence is only understood under sampling distributions that require potentially prohibitively expensive preprocessing steps. To address this limitation, we analyze RBK when the data is sampled uniformly, showing that its iterates converge in a Monte Carlo sense to a weighted least-squares solution. Unfortunately, for general problems the condition number of the weight matrix and the variance of the iterates can become arbitrarily large. We resolve these issues by incorporating regularization into the RBK iterations. Numerical experiments, including examples arising from natural gradient optimization, suggest that the regularized algorithm, ReBlocK, outperforms minibatch stochastic gradient descent for realistic problems that exhibit fast singular value decay.
Machine Learning, SciML, Natural Gradient Descent, Linear Regression, Kaczmarz
@@ 1 Introduction
$\begin{split}&\min_{x\in\mathbb{R}^{n}}\left\lVert Ax-b\right\rVert^{2},\\ A&\in\mathbb{R}^{m\times n},\;b\in\mathbb{R}^{m}.\end{split}$ (1)
@@
In this work, we are interested in algorithms that solve (1) by sampling just a small number of rows at a time. Such algorithms can be useful for extremely large problems for which direct methods and Krylov subspace methods are prohibitively expensive (Censor, 1981). Of particular interest are problems in which the dimension $n$ is so large that it is only possible to access $k\ll n$ rows at a time, but not so large that it is necessary to take $k=1$. As a canonical example, we might have $m=10^{9}$, $n=10^{6}$, and $k=10^{3}$.
As further motivation for algorithms of this type, there are some applications in which sampling rows is the only efficient way to access the data. For example, the rows may be computed on the fly, especially when solving the “semi-infinite” version of (1) in which the rows are indexed by continuous variables rather than discrete integers (Shustin & Avron, 2022). An example that captures both of these features is the problem of calculating natural gradient directions for continuous function learning problems; see Section 5.
One well-known approach for solving (1) using only a few rows at a time is minibatch stochastic gradient descent (mSGD). The mSGD algorithm works by sampling $k$ rows uniformly at random and using them to calculate an unbiased estimator of the gradient of the least-squares loss function. This is equivalent to averaging the $k$ independent gradient estimators furnished by each sampled row. For a thorough discussion of mSGD, see (Jain et al., 2018b).
The technique of averaging used in mSGD, while simple and efficient, is a relatively crude way of processing a block of $k$ rows. A more sophisticated approach is provided by the randomized block Kaczmarz method, which goes beyond averaging by making use of the pseudoinverse of the sampled block (Needell & Tropp, 2014). As we shall see, the use of the pseudoinverse opens up the possibility of faster convergence, but it also creates a number of complications regarding the implementation and analysis of the algorithm.
@@ 1.2 Randomized Kaczmarz
The modern version of the randomized Kaczmarz method (RK) was proposed by Thomas Strohmer and Roman Vershynin (2009). In RK, an initial guess $x_{0}$ is updated iteratively using a two-step procedure:
@@
    $x_{t+1}=x_{t}+a_{i_{t}}\frac{b_{i_{t}}-a_{i_{t}}^{\top}x_{t}}{\left\lVert a_{i_{t}}\right\rVert^{2}}.$ (3)
To reduce the sampling cost, it is also possible to run RK with uniform sampling, which is equivalent to running RK on a diagonally reweighted problem (Needell et al., 2014).
The iteration (3) has the interpretation of projecting $x_{t}$ onto the hyperplane of solutions to the individual equation $a_{i_{t}}^{\top}x=b_{i_{t}}$. For consistent systems, $Ax^{*}=b$, the RK iterates $x_{t}$ converge linearly to $x^{*}$ with a rate that depends on the conditioning of $A$. For inconsistent systems, $Ax^{*}\neq b$, RK converges only to within a finite horizon of the ordinary least-squares solution $x^{*}$ (Needell, 2010). In particular, the expected squared error $\mathbb{E}\,\left\lVert x_{t}-x^{*}\right\rVert^{2}$ converges to a finite, nonzero value that depends on the conditioning of $A$ and the norm of the residual vector $r=b-Ax^{*}$. See Theorem 7 of (Zouzias & Freris, 2013) for the strongest known convergence bound of this type.
@@ 1.3 Tail Averaging
Tail averaging is a common technique for boosting the accuracy of stochastic algorithms (Rakhlin et al., 2011; Jain et al., 2018b; Epperly et al., 2024). Given a series of stochastic iterates $x_{0},\ldots,x_{T}$ and a burn-in time $T_{b}$, the tail-averaged estimator is given by
$\overline{x}_{T}=\frac{1}{T-T_{b}}\sum_{t=T_{b}+1}^{T}x_{t}.$ (4)
The recent work (Epperly et al., 2024) shows that applying tail averaging to the RK iterates yields exact convergence (with no finite horizon) to the ordinary least-squares solution $x^{*}$, even for inconsistent systems. Building on these results, we will make use of tail averaging to obtain exact convergence to a weighted least-squares solution in the block case.
@@ 1.4 Randomized Block Kaczmarz
Randomized block Kaczmarz (RBK) is an extension of RK which uses blocks of rows to accelerate the convergence and make better use of parallel and distributed computing resources (Elfving, 1980; Needell & Tropp, 2014). Like RK, each RBK iteration proceeds in two steps:
@@
    $x_{t+1}=x_{t}+A_{S_{t}}^{+}(b_{S_{t}}-A_{S_{t}}x).$ (5)
Here $A_{S_{t}}^{+}$ is the Moore-Penrose pseudoinverse of $A_{S_{t}}$. The iteration (5) has the interpretation of projecting $x_{t}$ onto the hyperplane of solutions to the block of equations $A_{S_{t}}x=b_{S_{t}}$. The RBK method can also be viewed as a “sketch-and-project” algorithm; see (Gower & Richtárik, 2015).
There have been many proposals for how to choose the blocks in the RBK method. One idea is to use a preprocessing step to partition the matrix $A$ into well-conditioned blocks, then sample this fixed set of blocks uniformly (Needell & Tropp, 2014). It has also been suggested to preprocess the matrix with an incoherence transform, which can make it easier to generate a well-conditioned partition (Needell & Tropp, 2014) or, relatedly, enable RBK with uniform sampling to converge rapidly for the transformed problem (Dereziński & Yang, 2024). The work of (Dereziński & Yang, 2024) also provides an analysis of the RBK algorithm when sampling from a determinantal point process, and other proposals include greedy block Kaczmarz algorithms such as (Liu & Gu, 2021) which require evaluating the complete residual vector at each iteration.
Unfortunately many of these proposals apply only to consistent linear systems, and all of them require at least a preprocessing step in which the entire data matrix must be accessed. Such preprocessing can be prohibitively expensive for very large-scale problems, for which 1) it can be necessary to furnish an approximate solution without processing the entire data set even once (for instance, in the semi-infinite case), and 2) it can be impossible to manipulate more than a tiny subset of the data at a time due to storage constraints. This leads us to the central question of our work:
@@
$\min_{x\in\mathbb{R}^{n}}\mathbb{E}_{i\sim\mu}\left[(a_{i}^{\top}x-b_{i})^{2}\right],$ (6)
for which the corresponding algorithms would independently sample $k$ indices $i_{1},\ldots,i_{k}\sim\mu$. This is especially relevant for scientific applications, in which data is often continuous and may be sampled using a physics-based probability distribution. For example, in the case of neural network wavefunctions (Hermann et al., 2023), the sampling distribution is known as the Born probability density.
@@ 1.5 Contributions
Unfortunately, Theorem 3.1 does not guarantee that RBK-U is robust for every problem. To the contrary, when the problem contains many blocks $A_{S}$ that are nearly singular, the condition number of the weight matrix and the variance of the iterates can become arbitrarily large. See Figure 2 for numerical examples that manifest these issues.
One way to overcome these difficulties is to make stronger assumptions on the data. For example, Theorem 3.2 shows that convergence is obtained to the standard least-squares solution when the data arises from certain multivariate Gaussian distributions. Furthermore, in this case both the variance of the iterates and the convergence parameter $\alpha$ can be explicitly bounded. When the singular values of the covariance matrix decay rapidly, the convergence rate can be much faster than mSGD.
Many realistic problems are not well-modeled by Gaussian data. To provide a more general solution, we propose to regularize the RBK iterations as follows:
$x_{t+1}=x_{t}+A_{S_{t}}^{\top}(A_{S_{t}}A_{S_{t}}^{\top}+\lambda kI)^{-1}(b_{S_{t}}-A_{S_{t}}x_{t}).$ (7)
The RBK iteration (5) is recovered as $\lambda\rightarrow 0$, but we propose to instead use a small constant $\lambda>0$. This choice corresponds to a stochastic proximal point algorithm with a large, constant step size $\nicefrac{{1}}{{\lambda}}$, which to our knowledge has not been analyzed by previous works. We refer to this algorithm as the regularized block Kaczmarz method, or ReBlocK.
Similar to RBK-U, we show that ReBlocK with uniform sampling (ReBlocK-U) converges in a Monte Carlo sense to a weighted least-squares solution; see Theorem 4.1. Unlike for RBK-U, both the condition number of the weight matrix and the variance of the iterates can be controlled in terms of just $\lambda$ and some coarse properties of the data $A,b$. This makes ReBlocK-U much more reliable than RBK-U in practice. As an added benefit, ReBlocK iterations can be significantly more efficient than RBK iterations; see Sections 4.1 and F.2.1.
Our initial motivation for this work came from the problem of calculating natural gradient directions for deep neural networks. In Section 5, we explain how this setting naturally lends itself to the kinds of linear least-squares solvers we have studied in this paper. Encouragingly, Figure 3 shows that ReBlocK-U outperforms mSGD and RBK-U for this problem. Altogether, our results suggest that ReBlocK can be a more effective algorithm than mSGD for realistic problems that exhibit rapid singular value decay, especially when only moderate accuracy is required.
While our focus is on the case of uniform sampling, the same proof techniques can be applied to other sampling distributions. For example, in Appendix E we show that sampling from an appropriate determinantal point process can in theory enable tail averages of ReBlocK iterates to converge rapidly to the ordinary least-squares solution for inconsistent problems.
@@ 1.6 Related Works
Our regularized algorithm, ReBlocK, is closely related to the iterated Tikhonov-Kaczmarz method (De Cezaro et al., 2011). ReBlocK can also be viewed as a specific application of stochastic proximal point algorithms (sPPA) (Bertsekas, 2011; Asi & Duchi, 2019; Davis & Drusvyatskiy, 2019) for solving stochastic optimization problems with objective functions of the least-squares type. To ensure the exact convergence of sPPA, diminishing step sizes are required; see for example (Pătraşcu, 2021). In contrast, our work investigates the convergence of sPPA with a large constant step size for the special case of a least-squares loss. Such an approach allows for aggressive updates throughout the algorithm, potentially improving its practical efficiency.
Our work is also related to the nearly concurrent paper (Dereziński et al., 2025), which introduces Tikhonov regularization into the RBK iterations just like ReBlocK. (Dereziński et al., 2025) focuses on consistent systems that are preprocessed with a randomized Hadamard transform, and the regularization gives rise to optimal convergence rates in the presence of Nesterov acceleration. On the other hand, our work focuses on solving inconsistent systems without any preprocessing step, in which case the regularization is needed to ensure the stability of the algorithm.
The use of Nesterov acceleration represents a much broader trend in the development of stochastic iterative algorithms. Originally proposed by (Nesterov, 1983), this technique has been studied extensively in the context of both stochastic gradient and coordinate descent methods; see for example (Shalev-Shwartz & Zhang, 2013; Allen-Zhu et al., 2016; Jain et al., 2018a; Agarwal et al., 2020). Relative to block projection methods like RBK and ReBlocK, Nesterov acceleration represents an independent and complementary way to improve upon the convergence rate of mSGD. Incorporating Nesterov acceleration into our algorithms is a promising direction for future work.
@@ 2 Overview of the Analysis
$\overline{W}=\mathbb{E}_{S\sim\rho}\left[W(S)\right],\;\overline{P}=\mathbb{E}_{S\sim\rho}\left[P(S)\right].$ (11)
@@
$x^{(\rho)}=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\left\lVert Ax-b\right\rVert_{\overline{W}}^{2},\;r^{(\rho)}=b-Ax^{(\rho)}.$ (12)
@@
$x_{t+1}-x^{(\rho)}=(I-P(S_{t}))(x_{t}-x^{(\rho)})+A^{\top}W(S_{t})r^{(\rho)}.$ (13)
Since the normal equations for (12) can be written as $A^{\top}\overline{W}r^{(\rho)}=0$, we observe that the final term in (13) vanishes in expectation. Indeed, identifying the appropriate weighted solution $x^{(\rho)}$ to enable the generalized iteration (8) to be written in the form of (13), namely as a linear contraction of the error plus a zero-mean additive term, is the main technical innovation underlying our results. From here, the analysis of (Epperly et al., 2024) can be readily generalized to show that convergence to $x^{(\rho)}$ is obtained.

### v2

@@ Abstract
Due to the ever growing amounts of data leveraged for machine learning and scientific computing, it is increasingly important to develop algorithms that sample only a small portion of the data at a time. In the case of linear least-squares, the randomized block Kaczmarz method (RBK) is an appealing example of such an algorithm, but its convergence is only understood under sampling distributions that require potentially prohibitively expensive preprocessing steps. To address this limitation, we analyze RBK when the data is sampled uniformly, showing that its iterates converge in a Monte Carlo sense to a weighted least-squares solution. Unfortunately, for general problems the condition number of the weight matrix and the variance of the iterates can become arbitrarily large. We control these issues by incorporating regularization into the RBK iterations, yielding the regularized algorithm ReBlocK. Numerical experiments including examples arising from natural gradient optimization demonstrate that ReBlocK can outperform both RBK and minibatch stochastic gradient descent for inconsistent problems with rapidly decaying singular values.
keywords:
randomized Kaczmarz, linear least-squares, inconsistent problems, uniform sampling, regularization, natural gradient descent
{MSCcodes}
65F10, 65F20, 68W20, 15A06
@@ 1 Introduction
(1) $\begin{split}&\min_{x\in\mathbb{R}^{n}}\left\lVert Ax-b\right\rVert^{2},\\ A&\in\mathbb{R}^{m\times n},\;b\in\mathbb{R}^{m}.\end{split}$
@@
In this work, we are interested in algorithms that solve (1) by sampling just a small number of rows at a time. Such algorithms can be useful for extremely large problems for which direct methods and Krylov subspace methods are prohibitively expensive [6]. Of particular interest are problems in which the dimension $n$ is so large that it is only possible to access $k\ll n$ rows at a time, but not so large that it is necessary to take $k=1$. As a canonical example, we might have $m=10^{9}$, $n=10^{6}$, and $k=10^{3}$.
As further motivation for algorithms of this type, there are some applications in which sampling rows is the only efficient way to access the data. For example, the rows may be computed on the fly, especially when solving the “semi-infinite” version of (1) in which the rows are indexed by continuous variables rather than discrete integers [41]. An example that captures both of these features is the problem of calculating natural gradient directions for continuous function learning problems; see Section 5.
One well-known approach for solving (1) using only a few rows at a time is minibatch stochastic gradient descent (mSGD). The mSGD algorithm works by sampling $k$ rows uniformly at random and using them to calculate an unbiased estimator of the gradient of the least-squares loss function. This is equivalent to averaging the $k$ independent gradient estimators furnished by each sampled row. For a thorough discussion of mSGD, see [23].
The technique of averaging used in mSGD, while simple and efficient, is a relatively crude way of processing a block of $k$ rows. A more sophisticated approach is provided by the randomized block Kaczmarz method, which goes beyond averaging by making use of the pseudoinverse of the sampled block [31]. As we shall see, the use of the pseudoinverse opens up the possibility of faster convergence, but it also creates a number of complications regarding the implementation and analysis of the algorithm.
@@ 1.2 Randomized Kaczmarz
The modern version of the randomized Kaczmarz method (RK) was proposed by Thomas Strohmer and Roman Vershynin in 2009 [42]. In RK, an initial guess $x_{0}$ is updated iteratively using a two-step procedure:
@@
    (3) $x_{t+1}=x_{t}+a_{i_{t}}\frac{b_{i_{t}}-a_{i_{t}}^{\top}x_{t}}{\left\lVert a_{i_{t}}\right\rVert^{2}}.$
To reduce the sampling cost, it is also possible to run RK with uniform sampling, which is equivalent to running RK on a diagonally reweighted problem [32].
The iteration (3) has the interpretation of projecting $x_{t}$ onto the hyperplane of solutions to the individual equation $a_{i_{t}}^{\top}x=b_{i_{t}}$. For consistent systems, $Ax^{*}=b$, the RK iterates $x_{t}$ converge linearly to $x^{*}$ with a rate that depends on the conditioning of $A$. For inconsistent systems, $Ax^{*}\neq b$, RK converges only to within a finite horizon of the ordinary least-squares solution $x^{*}$ [30]. In particular, the expected squared error $\mathbb{E}\,\left\lVert x_{t}-x^{*}\right\rVert^{2}$ converges to a finite, nonzero value that depends on the conditioning of $A$ and the norm of the residual vector $r=b-Ax^{*}$. See Theorem 7 of [46] for the strongest known convergence bound of this type.
@@ 1.3 Tail Averaging
Tail averaging is a common technique for boosting the accuracy of stochastic algorithms [36, 23, 18]. Given a series of stochastic iterates $x_{0},\ldots,x_{T}$ and a burn-in time $T_{b}$, the tail-averaged estimator is given by
(4) $\overline{x}_{T}=\frac{1}{T-T_{b}}\sum_{t=T_{b}+1}^{T}x_{t}.$
The recent work [18] shows that applying tail averaging to the RK iterates yields exact convergence (with no finite horizon) to the ordinary least-squares solution $x^{*}$, even for inconsistent systems. Building on these results, we will make use of tail averaging to obtain exact convergence to a weighted least-squares solution in the block case.
@@ 1.4 Randomized Block Kaczmarz
Randomized block Kaczmarz (RBK) is an extension of RK which uses blocks of rows to accelerate the convergence and make better use of parallel and distributed computing resources [17, 31]. Like RK, each RBK iteration proceeds in two steps:
@@
    (5) $x_{t+1}=x_{t}+A_{S_{t}}^{+}(b_{S_{t}}-A_{S_{t}}x).$
Here $A_{S_{t}}^{+}$ is the Moore-Penrose pseudoinverse of $A_{S_{t}}$. The iteration (5) has the interpretation of projecting $x_{t}$ onto the hyperplane of solutions to the block of equations $A_{S_{t}}x=b_{S_{t}}$. The RBK method can also be viewed as a “sketch-and-project” algorithm; see [20].
There have been many proposals for how to choose the blocks in the RBK method. One idea is to use a preprocessing step to partition the matrix $A$ into well-conditioned blocks, then sample this fixed set of blocks uniformly [31]. It has also been suggested to preprocess the matrix with an incoherence transform, which can make it easier to generate a well-conditioned partition [31] or, relatedly, enable RBK with uniform sampling to converge rapidly for the transformed problem [15]. The work of [15] also provides an analysis of the RBK algorithm when sampling from a determinantal point process, and other proposals include greedy block Kaczmarz algorithms such as [26] which require evaluating the complete residual vector at each iteration.
Unfortunately most of these proposals apply only to consistent linear systems, and all of them require at least a preprocessing step in which the entire data matrix must be accessed. Such preprocessing can be prohibitively expensive for very large-scale problems, for which 1) it can be necessary to furnish an approximate solution without processing the entire data set even once (for instance, in the semi-infinite case), and 2) it can be impossible to manipulate more than a tiny subset of the data at a time due to storage constraints. This leads us to the central question of our work:
Question 1.1.
@@
(6) $\min_{x\in\mathbb{R}^{n}}\mathbb{E}_{i\sim\mu}\left[(a_{i}^{\top}x-b_{i})^{2}\right],$
for which the corresponding algorithms would independently sample $k$ indices $i_{1},\ldots,i_{k}\sim\mu$. This is especially relevant for scientific applications, in which data is often continuous and may be sampled using a physics-based probability distribution. For example, in the case of neural network wavefunctions [22], the sampling distribution is known as the Born probability density.
@@ 1.5 Contributions
In this work, we make substantial progress towards answering our central question 1.1. We first provide a new analysis of the RBK algorithm under uniform sampling which shows both the strengths and weaknesses of the algorithm. Inspired by this analysis, we propose and analyze a regularized algorithm, ReBlocK, which behaves more robustly for general types of data. Concretely, we summarize our contributions as follows:
  1. 1.
@@
  2. 2.
    We provide a new perspective on the pitfalls of RBK with uniform sampling. Concretely, our analysis reveals that when the problem contains many blocks $A_{S}$ that are nearly singular, the condition number of the weight matrix and the variance of the iterates can become arbitrarily large. In Fig. 3 we present numerical examples that manifest these issues.
  3. 3.
    We show that RBK-U performs robustly and efficiently for Gaussian data. Concretely, Theorem 3.2 shows that convergence is obtained to the standard least-squares solution when the data arises from a multivariate Gaussian distribution. Furthermore, in this case both the variance of the iterates and the convergence parameter $\alpha$ can be explicitly bounded. When the singular values of the covariance matrix decay rapidly, the convergence rate can be much faster than mSGD; see Corollary 3.3 and Fig. 1.
  4. 4.
    We propose and analyze a regularized algorithm to handle more general types of data. To provide a more general solution, we propose to regularize the RBK iterations as follows:
    (7) $x_{t+1}=x_{t}+A_{S_{t}}^{\top}(A_{S_{t}}A_{S_{t}}^{\top}+\lambda kI)^{-1}(b_{S_{t}}-A_{S_{t}}x_{t}).$
    We refer to this algorithm as the regularized block Kaczmarz method, or ReBlocK. Similar to RBK-U, we show that ReBlocK with uniform sampling (ReBlocK-U) converges in a Monte Carlo sense to a weighted least-squares solution; see Theorem 4.1. Unlike for RBK-U, both the condition number of the weight matrix and the variance of the iterates can be controlled in terms of just $\lambda$ and some coarse properties of the data $A,b$. This makes ReBlocK-U much more reliable than RBK-U in practice; see Fig. 3. As an added benefit, ReBlocK iterations can be significantly more efficient than RBK iterations; see Sections 4.1 and 8.
  5. 5.
    We demonstrate promising results for using ReBlocK to calculate natural gradient directions. Our initial motivation for this work came from the problem of calculating natural gradient directions for deep neural networks. In Section 5, we explain how this setting naturally lends itself to the kinds of linear least-squares solvers that we study in this paper. Encouragingly, Fig. 5 shows that ReBlocK-U outperforms mSGD and RBK-U when calculating natural gradient directions for a simple neural function regression task.
Altogether, our results demonstrate that regularization can improve the robustness of the randomized block Kaczmarz algorithm and unlock new applications to challenging problems for which sophisticated preprocessing schemes are out of reach. Furthermore, while our focus is on the case of uniform sampling, the same proof techniques can be applied to other sampling distributions. For example, in Appendix E we show that sampling from an appropriate determinantal point process can in theory enable tail averages of ReBlocK iterates to converge rapidly to the ordinary least-squares solution for arbitrary inconsistent problems.
@@ 1.6 Related Works
Our regularized algorithm, ReBlocK, is closely related to the iterated Tikhonov-Kaczmarz method [11]. ReBlocK can also be viewed as a specific application of stochastic proximal point algorithms (sPPA) [3, 2, 10] for solving stochastic optimization problems with objective functions of the least-squares type. To ensure the exact convergence of sPPA, diminishing step sizes are required; see for example [34]. In contrast, our work investigates the convergence of sPPA with a large constant step size for the special case of a least-squares loss. Such an approach allows for aggressive updates throughout the algorithm, potentially improving its practical efficiency.
The ReBlocK algorithm is additionally a special case of the SlimLS algorithm proposed by [8] for solving massive linear inverse problems. In fact, their convergence result for the expectation of the ReBlocK iterates, found in result (a) of their Theorem 3.1, is equivalent to equation (24) of our Theorem 4.1. However, by identifying the limiting expectation value as a weighted least-squares solution, we provide greater conceptual clarity regarding its meaning. Additionally, our bounds regarding the variance of the ReBlocK iterates, the convergence of tail averages, and the suboptimality of the weighted least-squares solution are all much stronger than the corresponding bounds of [8]. These stronger bounds serve to explain our numerical findings that ReBlocK can be robust and efficient even when $\lambda$ is very small and the problem contains many nearly singular blocks $A_{S}$, which is the most relevant case and is not explained by the previous analysis.
Our work is also related to the nearly concurrent paper [16], which introduces Tikhonov regularization into the RBK iterations just like ReBlocK. [16] focuses on consistent systems and in this context, the regularization gives rise to optimal convergence rates in the presence of Nesterov acceleration. On the other hand, our work focuses on solving inconsistent systems, in which case the regularization is needed to ensure the stability of the algorithm. Exploring the combination of regularization and Nesterov acceleration in the inconsistent case is a promising direction for future work.
@@ 2 Overview of the Analysis
(11) $\overline{W}=\mathbb{E}_{S\sim\rho}\left[W(S)\right],\;\overline{P}=\mathbb{E}_{S\sim\rho}\left[P(S)\right].$
@@
(12) $x^{(\rho)}=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\left\lVert Ax-b\right\rVert_{\overline{W}}^{2},\;r^{(\rho)}=b-Ax^{(\rho)}.$
@@
(13) $x_{t+1}-x^{(\rho)}=(I-P(S_{t}))(x_{t}-x^{(\rho)})+A^{\top}W(S_{t})r^{(\rho)}.$
Since the normal equations for (12) can be written as $A^{\top}\overline{W}r^{(\rho)}=0$, we observe that the final term in (13) vanishes in expectation. Indeed, identifying the appropriate weighted solution $x^{(\rho)}$ to enable the generalized iteration (8) to be written in the form of (13), namely as a linear contraction of the error plus a zero-mean additive term, is the main technical innovation underlying our results. From here, the analysis of [18] can be readily generalized to show that convergence to $x^{(\rho)}$ is obtained.